Tony makes deposits of $320 monthly (at the end of each month) into an account that pays 1.8% interest compounded monthly.

How much interest will he have earned after 3 years?


1
Expert's answer
2021-04-15T07:25:49-0400

"A=p(1+\frac{r}{(12)(100)})^{12\times n}=320(1+\frac{1.8}{(12)(100)})^{12\times 3}=320\times1.0554=\$337.728"

"I=A-P=337.728-320=\$17.728"